how to draw a rectangle with an area of 42 square units and a perimeter of 46 units help explain how you did it
Answer by mananth(16949)

let length be x
and width be y

Area = xy =42
y=42/x
Perimeter = 2(x+y)=46
x+y = 23
x+42/x = 23
multiply equation by x
x^2+42 = 23x
x^2-23x+42=0
x^2-21x-2x+42=0
x(x-21)-2(x-21)=0
(x-21)(x-2)=0
x=21 OR 2
The dimensions are 21 & 2 units